OpenStudy (anonymous):

Trignometric substitutio of the indefinite integral x^3*sqrt(16+x^2)dx

OpenStudy (owlfred):

Hoot! You just asked your first question! Hang tight while I find people to answer it for you. You can thank people who give you good answers by clicking the 'Good Answer' button on the right!

myininaya (myininaya):

just a sec k?

OpenStudy (anonymous):

k

myininaya (myininaya):

so far this is what i have

OpenStudy (anonymous):

kk looks good so far

myininaya (myininaya):

http://answers.yahoo.com/question/index?qid=20080508091701AAA0Mq0 im fixing to have dinner this link should help you with the rest im sorry

OpenStudy (anonymous):

the answer should be [(16-x^2)^5/2]/5-[16(16+x^2)^3/2]/3+C

myininaya (myininaya):

you should also put your answer back in terms of x i might be back on later if you haven't figured it out by then k?

OpenStudy (anonymous):

k thanks for your help.

myininaya (myininaya):

im sorry hai peace

OpenStudy (anonymous):

can you please help me?

OpenStudy (anonymous):

amistre its really tough

OpenStudy (amistre64):

x^3*sqrt(16+x^2) dx; yeah, it can be..

OpenStudy (amistre64):

4tan(t) = x^2

OpenStudy (amistre64):

ack!!.. 4tan(t) = x

OpenStudy (amistre64):

64 tan^3(t) * 4 sec(t) dt ; just gotta determine the dt now

OpenStudy (anonymous):

its 4sec^2 theta dtheta

OpenStudy (amistre64):

4sec^2(t) dt = dx right?

OpenStudy (anonymous):

yea

OpenStudy (amistre64):

64 tan^3(t) * 16sec^3(t) dt then right?

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

i put the answer above for reference, this problem is driving me crzy

OpenStudy (amistre64):

1024 (tan(t)sec(t))^3

OpenStudy (anonymous):

can you do it on your own and see if you get that?

OpenStudy (amistre64):

I aint got the time tonight, libraries closing soon but wolfram gives steps

OpenStudy (amistre64):

http://www.wolframalpha.com/input/?i=int%28x^3+sqrt%2816%2Bx^2%29%29dx or this for the original

OpenStudy (anonymous):

You've got it. You've worked your way to a function that can be integrated.

OpenStudy (anonymous):

Integral (sec theta tan theta)^3=sec^3 theta. Now you have to convert back. Good job Amistre. He just had to run and didn't have time to interpret it.